Twenty metrics of liquidity, Solvency, and Profitability

The comparative financial statements of Automotive Solutions Inc. are as follows. The market price of Automotive Solutions Inc. common stock was $64 on December 31, 20Y8.

AUTOMOTIVE SOLUTIONS INC.
Comparative Income Statement
For the Years Ended December 31, 20Y8 and 20Y7
    20Y8     20Y7
Sales $3,048,480 $2,808,690
Cost of goods sold (1,095,000) (1,007,400)
Gross profit $1,953,480 $1,801,290
Selling expenses $(702,710) $(838,590)
Administrative expenses (598,600) (492,510)
Total operating expenses (1,301,310) (1,331,100)
Operating income $652,170 $470,190
Other revenue and expense:
    Other income 34,330 30,010
    Other expense (interest) (216,000) (119,200)
Income before income tax $470,500 $381,000
Income tax expense (56,500) (45,900)
Net income $414,000 $335,100


AUTOMOTIVE SOLUTIONS INC.
Comparative Statement of Stockholders’ Equity
For the Years Ended December 31, 20Y8 and 20Y7
20Y8 20Y7
Preferred
Stock
Common
Stock
Retained
Earnings
Preferred
Stock
Common
Stock
Retained
Earnings
Balances, Jan. 1 $400,000 $460,000 $1,945,700 $400,000 $460,000 $1,636,000
Net income 414,000 335,100
Dividends:
    Preferred stock (7,000) (7,000)
    Common stock (18,400) (18,400)
Balances, Dec. 31 $400,000 $460,000 $2,334,300 $400,000 $460,000 $1,945,700


AUTOMOTIVE SOLUTIONS INC.
Comparative Balance Sheet
December 31, 20Y8 and 20Y7
    Dec. 31, 20Y8     Dec. 31, 20Y7
Assets
Current assets:
Cash $642,430 $360,500
Temporary investments 972,330 597,410
Accounts receivable (net) 540,200 511,000
Inventories 408,800 321,200
Prepaid expenses 121,549 72,100
Total current assets $2,685,309 $1,862,210
Long-term investments 1,052,721 6,760
Property, plant, and equipment (net) 2,970,000 2,673,000
Total assets $6,708,030 $4,541,970
Liabilities
Current liabilities $813,730 $246,270
Long-term liabilities:
Mortgage note payable, 8%, due in 15 years $1,210,000 $0
Bonds payable, 8%, due in 20 years 1,490,000 1,490,000
Total long-term liabilities $2,700,000 $1,490,000
Total liabilities $3,513,730 $1,736,270
Stockholders' Equity
Preferred $0.70 stock, $40 par $400,000 $400,000
Common stock, $10 par 460,000 460,000
Retained earnings 2,334,300 1,945,700
Total stockholders' equity $3,194,300 $2,805,700
Total liabilities and stockholders' equity $6,708,030 $4,541,970

Instructions:

Determine the following measures for 20Y8. Round ratio values to one decimal place and dollar amounts to the nearest cent. For number of days' sales in receivables and number of days' sales in inventory, round intermediate calculations to the nearest whole dollar and final amounts to one decimal place. Assume there are 365 days in the year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Days' sales in receivables days
6. Inventory turnover
7. Days' sales in inventory days
8. Debt ratio %
9. Ratio of liabilities to stockholders' equity
10. Ratio of fixed assets to long-term liabilities
11. Times interest earned times
12. Times preferred dividends earned times
13. Asset turnover
14. Return on total assets %
15. Return on stockholders’ equity %
16. Return on common stockholders’ equity %
17. Earnings per share on common stock $
18. Price-earnings ratio
19. Dividends per share of common stock $
20. Dividend yield %
Answer #1

1. Working capital = Current assets - Current Liabilities

= 2,685,309 - 813,730

= 1,871,579

2. Current ratio = Current assets / current liabilities

= 2,685,309/813,730

= 3.3

3. Quick Ratio = Quick assets / current liabilities

= (2,685,309-408,800-121,549)/813,730

= 2.6

4. Accounts Recievable turnover = Sales / Average accounts receivables

= 3,048,480 / (540,200+511,000)/2

= 3,048,480/ 525,600

= 5.8

As per HomeworkLib answering guidelines I answered first 4 sub parts of the question